OpenStudy (waheguru):

how do you solve this x+15=x/4

OpenStudy (waheguru):

please help fast

OpenStudy (anonymous):

\[4x+60=x \]

OpenStudy (anonymous):

4x+60=x 60=-3x x=-20

OpenStudy (waheguru):

negative how???

OpenStudy (anonymous):

im rit .. rit?

OpenStudy (anonymous):

\[3x= -60\]

OpenStudy (anonymous):

im rit ^^

OpenStudy (anonymous):

x= -20

OpenStudy (waheguru):

k thx pplz

OpenStudy (anonymous):

x+15=x/4 multiply by the LCD 4x+60=x

jhonyy9 (jhonyy9):

so 3x=-60 x= -20